OpenStudy (anonymous):

simplify: -16m^4y-4m^2y^6 / 8m^3y^4

OpenStudy (phi):

what do you get for the numbers -16 and 8 ?

OpenStudy (anonymous):

-2.

OpenStudy (phi):

yes. now do the m's m^4/m^3

OpenStudy (anonymous):

i got just m.

OpenStudy (phi):

yes.

OpenStudy (phi):

but I see another m^2 up top you can combine m*m^2 right ?

OpenStudy (anonymous):

so if i simplify that i get 1/m

OpenStudy (phi):

the problem is \[ \frac{ -16m^4y^{-4}m^2y^6}{8m^3y^4}\] you can replace -16/8 with -2 you can replace m^4/m^3 with m \[ \frac{ -2m y^{-4}m^2y^6}{y^4}\] notice up top you have m times m^2 (which you know is m*m*m) if you write that in short-hand you get m^3 up top

OpenStudy (anonymous):

alright i undertsand what step do i have to do now.

OpenStudy (phi):

you can also combine the y's up top y^-4 * y^6 up top is the same as y^6/y^4 (put the y in the bottom, and change -4 to +4) simplify to get y^2 now you still have y^4 in the bottom...

OpenStudy (anonymous):

so thats the answer or more steps.

OpenStudy (phi):

I have lost track of what is what. post what you have so far

OpenStudy (anonymous):

im kind of lost too

OpenStudy (phi):

start at the beginning, and write down the parts you have simplified, and leave the unsimplified parts (in any) as is.

OpenStudy (anonymous):

alright ill try even though im confused now lol

OpenStudy (phi):

the good news is you can simplify each letter individually... all the m's then all the y's

OpenStudy (anonymous):

my choices are 1. -5m^3y^3/2 2. 1/y 3. -2m/y^3 - y^2/2m 4. none 5. 2m/y^3+y^2/2m

OpenStudy (phi):

Based on those choices, I think you should use the equation editor to enter the original problem... Can you do that ?

OpenStudy (phi):

I think the problem is \[ \frac{-16m^4y-4m^2y^6}{ 8m^3y^4} \] you can write this as two separate fractions \[ \frac{-16m^4y}{ 8m^3y^4} -\frac{4m^2y^6}{ 8m^3y^4}\] now simplify each fraction